Already have an account? Get multiple benefits of using own account!
Login in your account..!
Remember me
Don't have an account? Create your account in less than a minutes,
Forgot password? how can I recover my password now!
Enter right registered email to receive password!
Marginal Probability
Probability of event A happening, denoted by P(A), is called single probability, marginal or unconditional probability.
Marginal or Unconditional Probability is defined as the ratio of number of possible outcomes favorable to the event A to the total number of possible outcomes.
P(A)
The definition assumes that the elements of the sample space have an equally likely chance of occurring.
Example
A gambler places a bet on numbers 14 through 25. There are 12 equally likely winning outcomes. The roulette wheel (a gambling instrument which can display any one of 38 equally likely numbers as the winning number) contains 38 equally likely outcomes.
The probability of the wheel stopping on a number from 14 through 25 (say event A) = 12/38 = 0.316.
The probability of losing, i.e. the wheel stopping on numbers other than 14 through 25 (say event B) is the probability of the complement of A occurring. The complement of an event A is defined as A' , where A' represents the non-occurrence of event A. So, the probability of A' (B) = 26/38 = 0.684.
P (A') = P(B) = 1 - P(A) because A and A' are the only possible events and they are mutually exclusive events of the sample of 38 equally likely outcomes. Thus, P(A) + P (A') = 1 and P(A and A' ) is 0.
if .77x + x = 8966.60, what is the value of x?
If 65% of the populations have black eyes, 25% have brown eyes and the remaining have blue eyes. What is the probability that a person selected at random has (i) Blue eyes (ii) Bro
Q. How to Add Mixed Numbers? Ans. If you have to add mixed numbers, you might try this method first: First rewrite the mixed number as a whole number plus a fracti
If tanA+sinA=m and m2-n2 = 4vmn, show that tanA-sinA=n
how do I change this ratio to a fraction
The function A(t) = 5(0.7)^t was used to define the amount A in milliliters of a drug in the bloodstream t hours after the drug was ingested. Determine algebraically the time it wi
Using the same mean and standard deviation as mean m = 20.1 and a standard deviation s = 5.8. Joe was informed that he scored at the 68 th percentile on the ACT, what was Joe's ap
as part of the markwting mix
INTRODUCTION : When a Class 5 child was given the problem 'If I paid Rs. 60 for 30 pencil boxes, how much did b pencil box cost?', he said it would be 60 x 30 = 1800. This
an insurance salesman sells policies to 5 men, all of identical age in good health. the probability that a man of this particular age will be alive 30 years hence is 2/3.Find the p
Get guaranteed satisfaction & time on delivery in every assignment order you paid with us! We ensure premium quality solution document along with free turntin report!
whatsapp: +91-977-207-8620
Phone: +91-977-207-8620
Email: [email protected]
All rights reserved! Copyrights ©2019-2020 ExpertsMind IT Educational Pvt Ltd